OpenStudy (staceyg):

Which of the following is equivalent to x^5y^2 / xy^2 when x /= 0 and y /= 0.

OpenStudy (staceyg):

x^6y^5 x^5y x^4y x^4

OpenStudy (anonymous):

Since \(x\not=0\), that means given the rational expression \(\dfrac{x}{x}\), you can simplify it to 1. So, \[\frac{x^5y^2}{xy^2}=x^4\] because you can eliminate the factors of \(y^2\) and one factor of \(x\).

OpenStudy (staceyg):

Thanks for the help :)